A marble and a cube are placed at the top of a ramp. Starting from rest at the same height, the marble rolls without slipping and the cube slides (no kinetic friction) down the ramp. Determine the ratio of the center-of-mass speed of the cube to the center-of-mass speed of the marble at the bottom of the ramp.
1
Expert's answer
2019-04-11T09:40:25-0400

The cube. From the conservation of energy:


"mgh = 0.5mv^2"

"v=\sqrt{2gh}"

The marble. From the conservation of energy:


"0.5MV^2 + 0.5 I\u03c9^2 = Mgh"

For a sphere:


"I=0.4 Mr^2"

"\omega=\frac{V}{r}"

Thus,


"0.5MV^2 + 0.5 (0.4 Mr^2)(\dfrac{V}{r})^2 = Mgh"

"V=\sqrt{\frac{10gh}{7}}"

So,


"\frac{v}{V}=\sqrt{\frac{7}{5}}"